the number of complaints fell from 50 per month to 35 per month after the training seminar. what was the percent of decrease?
Answer by ewatrrr(24785)


Hi
number of complaints fell from 50 per month to 35 per month after the training seminar.
what was the percent of decrease? (50-35)/50 = 15/50 = .30 OR 30%